OpenStudy (anonymous):

Write 15degrees26'9" using decimal degrees. Please walk me through it. Don't just give me the answer.

OpenStudy (anonymous):

\(26''\) means \(\frac{26}{60}\)

OpenStudy (anonymous):

oops that was wrong, sorry \(26'=\frac{26}{60}\)

OpenStudy (anonymous):

and \(p''=\frac{9}{3600}\) the denominator because \(60\times 60=3600\)

OpenStudy (anonymous):

weird use of base 60 instead of 10 because that is what the babylonians used as the base of their number system

OpenStudy (anonymous):

So would the answer be 15.437?

OpenStudy (anonymous):

yeah that looks good

OpenStudy (anonymous):

actually i am getting \(15.4358...\)

OpenStudy (anonymous):

Okay, that makes more sense. So when doing the opposite and writing 45.832degrees in degrees, minutes and seconds, would it be 45degrees49'45" or 45degrees49'55"?

OpenStudy (anonymous):

probably let me check

OpenStudy (anonymous):

yes, you are right

OpenStudy (anonymous):

Okay cool. Thanks!

OpenStudy (anonymous):

yw